OpenStudy (anonymous):

solve for x. 20x^2-7x-6=0 enter multiple answers by separating them by a comma.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

3/4, -2/5

OpenStudy (anonymous):

first factor \[(4x-3)(5x+2)=0\] etc then get kanwar's answers

OpenStudy (anonymous):

Kanwar's answers can be obtained as follows: 4x-3=0 => 4x=3 => x=3/4 5x-2=0 => 5x=-2 => x=-2/5